OpenStudy (anonymous):

Adding and Subtracting with unlike denominators. Find the LCD of these fractions: 4/5x^2y, 8/3xy, 12/13xy^5 a: 195xy b:195xy^5 c:195x^2y^5 d:195x^4y^7

OpenStudy (anonymous):

you need a 5, a 3 and a 13 also an \(x^2\) and a \(y^5\)

OpenStudy (anonymous):

just like with numbers each factor you see, to the highest power you see it in this case \[3\times 5\times 13x^2y^5\]

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

suppose i want the common denominator or least common multiple of \(18,24,30\)

OpenStudy (anonymous):

\[18=2\times 3^2\\ 24=2^3\times 3\\ 30=2\times 3\times 5\]

OpenStudy (anonymous):

my common multiple would be \[2^3\times 3^2\times 5=360\]

OpenStudy (anonymous):

you need each factor you see to the highest power you see it in any one term \[5x^2y\\ 3xy\\ 13xy^5\] you need all three numbers you need \(x^2\) and you need \(y^5\)

OpenStudy (anonymous):

so what does that mean